How many atoms are there in 5.10 moles of sulfur (S)? How many moles of calcium (Ca) atoms are in 77.4 g of Ca? Calculate the molar mass of a compound if 0.372 moles of it has a mass of 152 g.
Answer #1

No of atoms of S = n*avagadros number(Na)

                           = 5.1*6.023*10^23

                           = 3.072*10^24 atoms

No of mole of Ca = w/mwt = (77.4/40) = 1.935 mol

Molarmass of compound = w/n = 152/0.372 = 408.6 g/mol